equations in quadratic form;
solve the equation for x;
x^6-28x^3+27=0

Answer by rothauserc(4718)

x^6-28x^3+27=0
factor polynomial
(x^3 -27)(x^3-1) = 0
we have two cases
x^3 -27 = 0
x = 3
and
x^3 -1 = 0
x = 1
our two solutions are x = 3 or 1